Solve wquation by the zero-factor property
(w-1)(3w+2)=4w

Answer by dabanfield(803)

Solve wquation by the zero-factor property
(w-1)(3w+2)=4w
Expanding the above:
w*3w + 2w - 3w - 2 = 4w
3w^2 - 5w - 2 = 0
(3w + 1)*(w - 2) = 0
So we know that either;
3w + 1 = 0 or
w - 2 = 0.
Solve each of the above for w.